OpenStudy (anonymous):

2x + 5/2 > 3/2x -2 SOLVE FOR X

OpenStudy (anonymous):

would I add negative 2 to both sides? or make it a fraction too?

OpenStudy (anonymous):

That's one way to do it. If your plan is to eventually isolate the x on the right side. If you wanted to isolate the x on the left side you would subtract 5/2 from both sides.

OpenStudy (anonymous):

How would I solve and get what number X is supposed to be?

OpenStudy (anonymous):

sorry actually you want to add 2 to both sides not -2 First step 2x + 5/2 +2 > 3/2x -2 +2 2x + 9/2 > 3/2x

OpenStudy (anonymous):

Do you have an idea for what to do next?

OpenStudy (anonymous):

Our goal is to isolate the x on the right side, so we need all the x's on the right. How can I move that 2x over to the right?

OpenStudy (anonymous):

I would normally say combine like terms lol

OpenStudy (anonymous):

Yes, we want to move that 2x over to the right and combine it with the 3/2x. To move it we need to subtract 2x from both sides.

OpenStudy (anonymous):

but how would you do it with the fraction? BC idk how you got the 9/2

OpenStudy (anonymous):

5/2 + 2 = 5/2 + 4/2 = 9/2

OpenStudy (anonymous):

common denominator? is that how you got it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Can you write the next step after 2x + 9/2 > 3/2x

OpenStudy (anonymous):

9/2 > 7/2x ???

OpenStudy (anonymous):

So X would equal 7/2?

OpenStudy (anonymous):

Subtract 2x from both sides 2x + 9/2 -2x > 3/2x - 2x 9/2 > 3/2x - 4/2x 9/2 > -1/2x Your left side is good. Your right side is wrong

OpenStudy (anonymous):

I added the 2x

OpenStudy (anonymous):

Now, we just have to get rid of that -1/2 to isolate the x. How do we get rid of that -1/2?

OpenStudy (anonymous):

flip it?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

flip it and multiply by the flipped fraction on both sides. yes

OpenStudy (anonymous):

so -1/2 times -2/1?

OpenStudy (anonymous):

yes. that just equals 1.

OpenStudy (anonymous):

so x is 1?

OpenStudy (anonymous):

9/2 (-2/1) > -1/2x (-2/1) -18/2 > x -9 > x or x < -9 x is less than - 9 on inequalities you get x is < or > something . you don't get x = something

OpenStudy (anonymous):

I am so confused then

OpenStudy (anonymous):

general strategy. get all your x terms on one side and your numbers without x's on the other combine like terms. multiply both sides by the flipped version of your x coefficent

OpenStudy (anonymous):

You're confused by what I did in the last few steps?

OpenStudy (anonymous):

bc it says solve for x, I thought I was supposed to get a number for x

OpenStudy (anonymous):

SO 9/2 > X WOULD BE THE CORRECT ANSWER?

OpenStudy (anonymous):

you did solve for x. with inequalities x does not equal just one number though. x can be any number less than -9. Pick any number less than -9 and your original inequality, 2x + 5/2 > 3/2x -2 will be true.

OpenStudy (anonymous):

No, the answer is x < -9

OpenStudy (anonymous):

see, idk how you flipped it and got just the -9

OpenStudy (anonymous):

ooook, I saw your post just now, I didn't see it earlier about flipping it

OpenStudy (anonymous):

you have to do the same thing to both sides in algebra. golden rule there. I had to multiply the right side by -2/1 also. go back up and look. I multiplied 9/2 by -2/1 multiply numerators and denominators to get -18/2 which simplifies to -9/1 or -9

OpenStudy (anonymous):

I have one last problem.

OpenStudy (anonymous):

uhoh

OpenStudy (anonymous):

-3 ( x+1) - x/2 + 3/2 < 0

OpenStudy (anonymous):

I distribute the -3 to whats in the ( ) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so far I have -3x -3 - x/2 + 3/2 < 0

OpenStudy (anonymous):

yes. now combine those x terms (that x/2 is the same as (1/2)x) and combine the -3 and +3/2

OpenStudy (anonymous):

would that be - 4x/2?

OpenStudy (anonymous):

bc I found the common denominators

OpenStudy (anonymous):

no. -3x - (1/2)x = -6/2x - (1/2)x = (-7/2)x

OpenStudy (anonymous):

ok, I see where I went wrong.

OpenStudy (anonymous):

so the next thing would be to add -3 to 3/2 and I got 3/2 for the answer?

OpenStudy (anonymous):

no. -3 + 3/2 = -6/2 + 3/2 = -3/2 that minus sign is important

OpenStudy (anonymous):

So I have the 3/2 right, just not the negative right?

OpenStudy (anonymous):

yes. so what do we have now? (-7/2)x - 3/2 < 0 what's next?

OpenStudy (anonymous):

adding the -3/2 to both sides?

OpenStudy (anonymous):

add 3/2 to both sides

OpenStudy (anonymous):

so I was right?

OpenStudy (anonymous):

except for the -...yes

OpenStudy (anonymous):

(-7/2)x - 3/2 + 3/2 < 0 + 3/2 (-7/2)x < 3/2 Now multiply both sides by -2/7 to get x < 3/2 (-2/7) x < -6/14 x < -3/7 sorry gotta run. good luck!

OpenStudy (anonymous):

thank you